# A Variational Principle for Extended Irreversible Thermodynamics: Heat Conducting Viscous Fluids

### v1

@@ 3.6.1 Entropy flux beyond local equilibrium
Variational principle and evolution equations.
Since only the flux of entropy is changed, it suffices to add in our variational principle an internal entropy source in the form of a flux, denoted as $\mathcal{J}=\operatorname{div}\boldsymbol{J}^{\prime}$, where $\boldsymbol{J}^{\prime}$ represents the part of the entropy flux not accounted for in the heat conduction process.
@@
$\left\{\!\!\begin{array}[]{l}\vspace{0.2cm}\displaystyle\frac{d}{dt}\frac{\partial\widehat{\mathfrak{L}}}{\partial\dot{\boldsymbol{\varphi}}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{\varphi}}=\operatorname{DIV}\Big{(}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\nabla{\boldsymbol{\varphi}}}+\boldsymbol{P}\Big{)}\\ \displaystyle-\frac{\partial\widehat{\mathfrak{L}}}{\partial S}(\dot{S}+\operatorname{DIV}(\boldsymbol{J}+\boldsymbol{J}^{\prime}))-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}:\dot{\boldsymbol{S}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\cdot\dot{\boldsymbol{Q}}=\boldsymbol{S}:\boldsymbol{D}(\dot{\boldsymbol{\varphi}})+\boldsymbol{J}\cdot\nabla\frac{\partial\widehat{\mathfrak{L}}}{\partial S}-\frac{\partial\widehat{\mathfrak{L}}}{\partial S}\operatorname{DIV}\boldsymbol{J}^{\prime},\end{array}\right.$ (3.56)
@@
^$\frac{\partial\widehat{\ell}}{\partial s}\bar{D}_{t}\varsigma+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}=-\boldsymbol{\sigma}:\nabla\boldsymbol{u}+\boldsymbol{j}\cdot\nabla D_{t}\gamma-jD_{t}\gamma,$ (3.58)
@@
^$\frac{\partial\widehat{\ell}}{\partial s}\bar{D}_{\delta}\varsigma+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{\delta}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{\delta}\boldsymbol{q}=-\boldsymbol{\sigma}:\nabla\boldsymbol{\zeta}+\boldsymbol{j}\cdot\nabla D_{\delta}\gamma-jD_{\delta}\gamma$ (3.59)
@@
$\left\{\begin{array}[]{l}\vspace{0.2cm}\displaystyle\partial_{t}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}+\pounds_{\boldsymbol{u}}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}=\rho\nabla\frac{\partial\widehat{\ell}}{\partial\rho}+s\nabla\frac{\partial\widehat{\ell}}{\partial s}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\nabla\boldsymbol{\sigma}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\nabla\boldsymbol{q}+\operatorname{div}(\boldsymbol{\sigma}+\boldsymbol{\tau}_{\boldsymbol{\sigma}}+\boldsymbol{\tau}_{\boldsymbol{q}})\\ \vspace{0.2cm}\displaystyle-\frac{\partial\widehat{\ell}}{\partial s}(\bar{D}_{t}s+\operatorname{div}(\boldsymbol{j}+\boldsymbol{j}^{\prime}))=\boldsymbol{\sigma}:\nabla\boldsymbol{u}+\boldsymbol{j}\cdot\nabla\frac{\partial\widehat{\ell}}{\partial s}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}-\frac{\partial\widehat{\ell}}{\partial s}\operatorname{div}\boldsymbol{j}^{\prime}\\ \displaystyle\bar{D}_{t}\rho=0,\end{array}\right.$ (3.60)
@@
$\left\{\begin{array}[]{l}\vspace{0.2cm}\displaystyle\rho(\partial_{t}\boldsymbol{u}+\boldsymbol{u}\cdot\nabla\boldsymbol{u})=-\nabla p+\operatorname{div}(\boldsymbol{\tau}_{\boldsymbol{q}}+\boldsymbol{\tau}_{\boldsymbol{\sigma}}+\boldsymbol{\sigma})\\ \vspace{0.2cm}\displaystyle T(\bar{D}_{t}s+\operatorname{div}(\boldsymbol{j}+\boldsymbol{j}^{\prime})=\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T-\frac{\partial\widehat{\varepsilon}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}-\frac{\partial\widehat{\varepsilon}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}+T\operatorname{div}\boldsymbol{j}^{\prime}\\ \displaystyle\bar{D}_{t}\rho=0.\end{array}\right.$ (3.62)
Balance laws and thermodynamic consistency.
We note that this modification doesn’t affect the energy balance derived earlier, as $\boldsymbol{j}^{\prime}$ simply does not appear in it. The entropy balance (3.47) is however changed into
@@
$\frac{d}{dt}\int_{\mathcal{D}}\!s\,{\rm d}\mathbf{x}=\int_{\mathcal{D}}\!\frac{1}{T}\bigg{(}\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}-\frac{\partial\widehat{\ell}}{\partial s}\operatorname{div}\boldsymbol{j}^{\prime}\bigg{)}{\rm d}\mathbf{x}-\int_{\partial\mathcal{D}}\!\boldsymbol{j}_{s}\cdot\boldsymbol{n}\,{\rm d}A.$
@@
$\boldsymbol{\sigma}:\big{(}\operatorname{Def}\boldsymbol{u}-\beta\mathfrak{D}_{t}\boldsymbol{\sigma}+\mathbb{M}\cdot\nabla\boldsymbol{q}\big{)}+\boldsymbol{q}\cdot\big{(}-\frac{1}{T}\nabla T-\alpha\mathfrak{D}_{t}\boldsymbol{q}+\nabla\boldsymbol{\sigma}\cdot\mathbb{M}\big{)}\geq 0.$
@@
$\boldsymbol{\sigma}=\mathbb{L}\big{(}\operatorname{Def}\boldsymbol{u}-\beta\mathfrak{D}_{t}\boldsymbol{\sigma}+\mathbb{M}\cdot\nabla\boldsymbol{q}\big{)},\qquad\boldsymbol{q}=\mathbb{K}\big{(}-\frac{1}{T}\nabla T-\alpha\mathfrak{D}_{t}\boldsymbol{q}+\nabla\boldsymbol{\sigma}\cdot\mathbb{M}\big{)}$
@@ 3.6.2 Higher-order fluxes
Variational principle and evolution equations.
In this case, the nonequilibrium Lagrangian density in the material description takes the form $\mathfrak{L}(\boldsymbol{\varphi},\dot{\boldsymbol{\varphi}},\nabla\boldsymbol{\varphi},\varrho,S,\boldsymbol{Q}^{(1)},...,\boldsymbol{Q}^{(n)})$, where $\boldsymbol{Q}^{(k)}$ is a $k$-times contravariant tensor field density. The higher-order fluxes are naturally incorporated in the phenomenological constraint (3.54) as
@@
Similarly, in the Eulerian description, the variational condition (3.28) remains unchanged, except that it now involves a nonequilibrium Lagrangian density ^$\widehat{\ell}(\boldsymbol{u},\rho,s,\boldsymbol{q}^{(1)},...,\boldsymbol{q}^{(n)})$, which depends on the Eulerian higher-order fluxes $\boldsymbol{q}^{(k)}=\boldsymbol{\varphi}_{*}\boldsymbol{Q}^{(k)}J_{\boldsymbol{\varphi}^{-1}}$. Naturally, this requires the Lagrangian density $\mathfrak{L}$ to be material covariant, see (3.13). The phenomenological constraint (3.58) is modified to
^$\frac{\partial\widehat{\ell}}{\partial s}\bar{D}_{t}\varsigma+\sum_{k=1}^{n}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}^{(k)}}:\mathfrak{D}_{t}\boldsymbol{q}^{(k)}=\boldsymbol{j}\cdot\nabla D_{t}\gamma-jD_{t}\gamma,$ (3.66)
@@
^$\frac{\partial\widehat{\ell}}{\partial s}\bar{D}_{\delta}\varsigma+\sum_{k=1}^{n}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}^{(k)}}:\mathfrak{D}_{\delta}\boldsymbol{q}^{(k)}=\boldsymbol{j}\cdot\nabla D_{\delta}\gamma-jD_{\delta}\gamma$ (3.67)
which replaces (3.59). They incorporate the rates $\mathfrak{D}_{t}$ and variations $\mathfrak{D}_{\delta}$ associated with the type of tensorial density for $k=1,...,n$, and follow from (3.64) and (3.65) via the Lagrange-to-Euler transformation, as before. The resulting system is
$\left\{\begin{array}[]{l}\vspace{0.2cm}\displaystyle\partial_{t}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}+\pounds_{\boldsymbol{u}}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}=\rho\nabla\frac{\partial\widehat{\ell}}{\partial\rho}+s\nabla\frac{\partial\widehat{\ell}}{\partial s}-\sum_{k=1}^{n}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}^{(k)}}:\nabla\boldsymbol{q}^{(k)}+\operatorname{div}\Big{(}\sum_{k=1}^{n}\boldsymbol{\tau}_{\boldsymbol{q}^{(k)}}\Big{)}\\ \vspace{0.2cm}\displaystyle-\frac{\partial\widehat{\ell}}{\partial s}(\bar{D}_{t}s+\operatorname{div}(\boldsymbol{j}+\boldsymbol{j}^{\prime}))=\boldsymbol{j}\cdot\nabla\frac{\partial\widehat{\ell}}{\partial s}+\sum_{k=1}^{n}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}^{(k)}}:\mathfrak{D}_{t}\boldsymbol{q}^{(k)}-\frac{\partial\widehat{\ell}}{\partial s}\operatorname{div}\boldsymbol{j}^{\prime}\\ \displaystyle\bar{D}_{t}\rho=0,\end{array}\right.$ (3.68)
@@
Balance laws and thermodynamic consistency.
The energy balance is the generalization of (3.4) to higher-order fluxes, while the entropy balance extends (3.63) as follows:
@@
$\frac{\partial\widehat{\varepsilon}}{\partial\boldsymbol{q}^{(k)}}=\alpha_{k}\boldsymbol{q}^{(k)},\qquad k=1,...,n.$
@@
$\sum_{k=1}^{n}\boldsymbol{q}^{(k)}\cdot\Big{(}-\frac{\nabla T}{T}\delta^{k}_{1}-\alpha_{k}\mathfrak{D}_{t}\boldsymbol{q}^{(k)}+\gamma_{k}T\operatorname{div}(\boldsymbol{q}^{(k+1)})+T\gamma_{k-1}\nabla\boldsymbol{q}^{(k-1)}\Big{)}\geq 0,$
@@
$\tau_{k}\big{(}\mathfrak{D}_{t}\boldsymbol{q}^{(k)}+\operatorname{div}(\boldsymbol{q}^{(k+1)})\big{)}+\boldsymbol{q}^{(k)}=-\kappa_{k}\nabla\boldsymbol{q}^{(k-1)},\quad k=1,...,n,$

### v2

@@ 3.6.1 Entropy flux beyond local equilibrium
Variational principle and evolution equations. Since only the flux of entropy is changed, it suffices to add in our variational principle an internal entropy source in the form of a flux, denoted as $\mathcal{J}=\operatorname{div}\boldsymbol{J}^{\prime}$, where $\boldsymbol{J}^{\prime}$ represents the part of the entropy flux not accounted for in the heat conduction process.
@@
$\hskip-8.5359pt\left\{\!\!\begin{array}[]{l}\vskip 5.69046pt\displaystyle\frac{d}{dt}\frac{\partial\widehat{\mathfrak{L}}}{\partial\dot{\boldsymbol{\varphi}}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{\varphi}}=\operatorname{DIV}\Big{(}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\nabla{\boldsymbol{\varphi}}}+\boldsymbol{P}\Big{)}\\ \displaystyle-\frac{\partial\widehat{\mathfrak{L}}}{\partial S}(\dot{S}+\operatorname{DIV}(\boldsymbol{J}+\boldsymbol{J}^{\prime}))-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}:\dot{\boldsymbol{S}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\cdot\dot{\boldsymbol{Q}}=\boldsymbol{S}:\boldsymbol{D}(\dot{\boldsymbol{\varphi}})+\boldsymbol{J}\cdot\nabla\frac{\partial\widehat{\mathfrak{L}}}{\partial S}-\frac{\partial\widehat{\mathfrak{L}}}{\partial S}\operatorname{DIV}\boldsymbol{J}^{\prime},\hskip-8.5359pt\end{array}\right$. (3.83)
@@
$\frac{\partial\widehat{\ell}}{\partial s}\bar{D}_{t}\varsigma+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}=-\boldsymbol{\sigma}:\nabla\boldsymbol{u}+\boldsymbol{j}\cdot\nabla D_{t}\gamma-jD_{t}\gamma,$ (3.85)
@@
$\frac{\partial\widehat{\ell}}{\partial s}\bar{D}_{\delta}\varsigma+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{\delta}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{\delta}\boldsymbol{q}=-\boldsymbol{\sigma}:\nabla\boldsymbol{\zeta}+\boldsymbol{j}\cdot\nabla D_{\delta}\gamma-jD_{\delta}\gamma$ (3.86)
@@
$\hskip-14.22636pt\left\{\begin{array}[]{l}\vskip 5.69046pt\displaystyle\partial_{t}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}+\pounds_{\boldsymbol{u}}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}=\rho\nabla\frac{\partial\widehat{\ell}}{\partial\rho}+s\nabla\frac{\partial\widehat{\ell}}{\partial s}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\nabla\boldsymbol{\sigma}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\nabla\boldsymbol{q}+\operatorname{div}(\boldsymbol{\sigma}+\boldsymbol{\tau}_{\boldsymbol{\sigma}}+\boldsymbol{\tau}_{\boldsymbol{q}})\\ \vskip 5.69046pt\displaystyle-\frac{\partial\widehat{\ell}}{\partial s}(\bar{D}_{t}s+\operatorname{div}(\boldsymbol{j}+\boldsymbol{j}^{\prime}))=\boldsymbol{\sigma}:\nabla\boldsymbol{u}+\boldsymbol{j}\cdot\nabla\frac{\partial\widehat{\ell}}{\partial s}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}-\frac{\partial\widehat{\ell}}{\partial s}\operatorname{div}\boldsymbol{j}^{\prime}\\ \displaystyle\bar{D}_{t}\rho=0,\hskip-8.5359pt\end{array}\right$. (3.87)
@@
$\left\{\begin{array}[]{l}\vskip 5.69046pt\displaystyle\rho(\partial_{t}\boldsymbol{u}+\boldsymbol{u}\cdot\nabla\boldsymbol{u})=-\nabla p+\operatorname{div}(\boldsymbol{\tau}_{\boldsymbol{q}}+\boldsymbol{\tau}_{\boldsymbol{\sigma}}+\boldsymbol{\sigma})\\ \vskip 5.69046pt\displaystyle T(\bar{D}_{t}s+\operatorname{div}(\boldsymbol{j}+\boldsymbol{j}^{\prime})=\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T-\frac{\partial\widehat{\varepsilon}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}-\frac{\partial\widehat{\varepsilon}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}+T\operatorname{div}\boldsymbol{j}^{\prime}\\ \displaystyle\bar{D}_{t}\rho=0.\end{array}\right$. (3.89)
Balance laws and thermodynamic consistency. We note that this modification doesn’t affect the energy balance derived earlier, as $\boldsymbol{j}^{\prime}$ simply does not appear in it. The entropy balance (3.66) is however changed into
@@
$\frac{d}{dt}\int_{\mathcal{D}}\!s\,{\rm d}\mathbf{x}=\int_{\mathcal{D}}\!\frac{1}{T}\bigg{(}\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}-\frac{\partial\widehat{\ell}}{\partial s}\operatorname{div}\boldsymbol{j}^{\prime}\bigg{)}{\rm d}\mathbf{x}-\int_{\partial\mathcal{D}}\!\boldsymbol{j}_{s}\cdot\boldsymbol{n}\,{\rm d}A.$ (3.91)
@@
$\boldsymbol{\sigma}:\big{(}\operatorname{Def}\boldsymbol{u}-\beta\mathfrak{D}_{t}\boldsymbol{\sigma}+\mathbb{M}\cdot\nabla\boldsymbol{q}\big{)}+\boldsymbol{q}\cdot\big{(}-\frac{1}{T}\nabla T-\alpha\mathfrak{D}_{t}\boldsymbol{q}+\nabla\boldsymbol{\sigma}\cdot\mathbb{M}\big{)}\geq 0.$ (3.92)
@@
$\boldsymbol{\sigma}=\mathbb{L}\big{(}\operatorname{Def}\boldsymbol{u}-\beta\mathfrak{D}_{t}\boldsymbol{\sigma}+\mathbb{M}\cdot\nabla\boldsymbol{q}\big{)},\qquad\boldsymbol{q}=\mathbb{K}\big{(}-\frac{1}{T}\nabla T-\alpha\mathfrak{D}_{t}\boldsymbol{q}+\nabla\boldsymbol{\sigma}\cdot\mathbb{M}\big{)}$ (3.93)
@@ 3.6.2 Higher-order fluxes
Variational principle and evolution equations. In this case, the nonequilibrium Lagrangian density in the material description takes the form $\mathfrak{L}(\boldsymbol{\varphi},\dot{\boldsymbol{\varphi}},\nabla\boldsymbol{\varphi},\varrho,S,\boldsymbol{Q}^{(1)},...,\boldsymbol{Q}^{(n)})$, where $\boldsymbol{Q}^{(k)}$ is a $k$-times contravariant tensor field density. The higher-order fluxes are naturally incorporated in the phenomenological constraint (3.81) as
@@
Similarly, in the Eulerian description, the variational condition (3.40) remains unchanged, except that it now involves a nonequilibrium Lagrangian density $\widehat{\ell}(\boldsymbol{u},\rho,s,\boldsymbol{q}^{(1)},...,\boldsymbol{q}^{(n)})$, which depends on the Eulerian higher-order fluxes $\boldsymbol{q}^{(k)}=\boldsymbol{\varphi}_{*}\boldsymbol{Q}^{(k)}J_{\boldsymbol{\varphi}^{-1}}$. Naturally, this requires the Lagrangian density $\mathfrak{L}$ to be material covariant, see (3.18). The phenomenological constraint (3.85) is modified to
$\frac{\partial\widehat{\ell}}{\partial s}\bar{D}_{t}\varsigma+\sum_{k=1}^{n}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}^{(k)}}:\mathfrak{D}_{t}\boldsymbol{q}^{(k)}=\boldsymbol{j}\cdot\nabla D_{t}\gamma-jD_{t}\gamma,$ (3.96)
@@
$\frac{\partial\widehat{\ell}}{\partial s}\bar{D}_{\delta}\varsigma+\sum_{k=1}^{n}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}^{(k)}}:\mathfrak{D}_{\delta}\boldsymbol{q}^{(k)}=\boldsymbol{j}\cdot\nabla D_{\delta}\gamma-jD_{\delta}\gamma$ (3.97)
which replaces (3.86). They incorporate the rates $\mathfrak{D}_{t}$ and variations $\mathfrak{D}_{\delta}$ associated with the type of tensorial density for k=1,,nk=1,...,n, and follow from (3.94) and (3.95) via the Lagrange-to-Euler transformation, as before. The resulting system is
$\hskip-14.22636pt\left\{\begin{array}[]{l}\vskip 5.69046pt\displaystyle\partial_{t}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}+\pounds_{\boldsymbol{u}}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}=\rho\nabla\frac{\partial\widehat{\ell}}{\partial\rho}+s\nabla\frac{\partial\widehat{\ell}}{\partial s}-\sum_{k=1}^{n}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}^{(k)}}:\nabla\boldsymbol{q}^{(k)}+\operatorname{div}\Big{(}\sum_{k=1}^{n}\boldsymbol{\tau}_{\boldsymbol{q}^{(k)}}\Big{)}\\ \vskip 5.69046pt\displaystyle-\frac{\partial\widehat{\ell}}{\partial s}(\bar{D}_{t}s+\operatorname{div}(\boldsymbol{j}+\boldsymbol{j}^{\prime}))=\boldsymbol{j}\cdot\nabla\frac{\partial\widehat{\ell}}{\partial s}+\sum_{k=1}^{n}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}^{(k)}}:\mathfrak{D}_{t}\boldsymbol{q}^{(k)}-\frac{\partial\widehat{\ell}}{\partial s}\operatorname{div}\boldsymbol{j}^{\prime}\\ \displaystyle\bar{D}_{t}\rho=0,\hskip-8.5359pt\end{array}\right$. (3.98)
@@
Balance laws and thermodynamic consistency. The energy balance is the generalization of (3.4) to higher-order fluxes, while the entropy balance extends (3.90) as follows:
@@
$\frac{\partial\widehat{\varepsilon}}{\partial\boldsymbol{q}^{(k)}}=\alpha_{k}\boldsymbol{q}^{(k)},\qquad k=1,...,n.$ (3.101)
@@
$\sum_{k=1}^{n}\boldsymbol{q}^{(k)}\cdot\Big{(}-\frac{\nabla T}{T}\delta^{k}_{1}-\alpha_{k}\mathfrak{D}_{t}\boldsymbol{q}^{(k)}+\gamma_{k}T\operatorname{div}(\boldsymbol{q}^{(k+1)})+T\gamma_{k-1}\nabla\boldsymbol{q}^{(k-1)}\Big{)}\geq 0,$ (3.103)
@@
$\tau_{k}\big{(}\mathfrak{D}_{t}\boldsymbol{q}^{(k)}+\operatorname{div}(\boldsymbol{q}^{(k+1)})\big{)}+\boldsymbol{q}^{(k)}=-\kappa_{k}\nabla\boldsymbol{q}^{(k-1)},\quad k=1,...,n,$ (3.104)